OpenStudy (anonymous):

2x/4x+2 * 14x+7/6 a. 7x/12 b. 7/6 c. x/6 d. 7x/6

OpenStudy (anonymous):

\[\frac{ 2x }{4x+2} * \frac{ 14x}{\frac{ 7 }{ 6 }}\] is that it?

OpenStudy (anonymous):

No let me draw it @jperez570

OpenStudy (anonymous):

oh wait \[14x+7/6\]

OpenStudy (anonymous):